OpenStudy (anonymous):

What is the range of the quadratic equation y= 2x^2 -8x +6? 1) range= y≥-8 2) range= y≤-8 3) range= y≤2 4) range= y≥2

OpenStudy (anonymous):

find the second coordinate of the vertex. first coordinate is \[-\frac{b}{2a}=-\frac{-8}{2\times 2}=2\]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

range=y>= -2

OpenStudy (anonymous):

second coordinate is 8 - 16 + 6 = -2

OpenStudy (anonymous):

\[y \ge -2\]\[[-2, \infty]\]

myininaya (myininaya):

y=2x^2-8x+6=2(x^2-4x)+6=2(x^2-4x+4)+6-2(4) =2(x-2)^2+6-8=2(x-2)^2-2 vertex is (2,-2) parabola is concave up [-2,inf)